OpenStudy (toshabay16):

Medal and fan please help 1. f(x) = x2 – 4 and g(x) = x – 2 (a) What is (f + g)(x)? (b) What is (f – g)(x)? (c) What is (fg)(x)? (d) What is ?

OpenStudy (toshabay16):

Medal and fan

OpenStudy (anonymous):

do you know what mean (f+g)(x) ?

OpenStudy (toshabay16):

No can you explain the problem to me and help me solve it

OpenStudy (photon336):

I believe you would for b subtract f(x)-g(x) a. add f(x)+g(x) c multiply the two functions together

OpenStudy (toshabay16):

how would i do that

OpenStudy (photon336):

f(x)+g(x) \[x^{2}-4+(x-2) = ? \]

OpenStudy (toshabay16):

ok

OpenStudy (toshabay16):

x^2+x-6?

Vocaloid (vocaloid):

yes

Vocaloid (vocaloid):

now, for b you just subtract f(x) - g(x) so x2 – 4 - ( x – 2 )

OpenStudy (toshabay16):

B)x2−3x−4

Vocaloid (vocaloid):

check your math

Vocaloid (vocaloid):

x^2 – 4 - ( x – 2 ) = x^2 - 4 - x + 2 = ?

OpenStudy (toshabay16):

x2−x−2

Vocaloid (vocaloid):

good now for part c) you just multiply (x2 – 4)(x – 2)

OpenStudy (toshabay16):

−x3−4x+8 @Vocaloid

OpenStudy (toshabay16):

x3-2x2-4x+8 @Vocaloid